OpenStudy (anonymous):

What is the simplified form of (b/2)/(4/2b)

OpenStudy (anonymous):

i got 4/b^2 is that right?

OpenStudy (anonymous):

flip the second one and multiply instead.....

OpenStudy (anonymous):

\[\frac{\frac{b}{2}}{\frac{4}{2b}} = \frac{b}{2} \times \frac{2b}{4}\]

OpenStudy (anonymous):

(b/2)/(4/2b) =(b/2) x (2b/4) = (b/2) x (b/2) =(b^2)/4

OpenStudy (hba):

hmm good